ITEM QUANTITY (2009) PRICE (2009) QUANTITY (2010) PRICE (2009) Bananas 20 $1.00 15 $1.00 Cupcakes 30...

ITEM

QUANTITY (2009)

PRICE (2009)

QUANTITY (2010)

PRICE (2009)

Bananas

20

$1.00

15

$1.00

Cupcakes

30

$1.00

45

$0.75

Sushi

10

$10.00

8

$11.00

  1. A Consumer Expenditure Survey in Fancyville shows that people only consume bananas, cupcakes and sushi. The Consumer Expenditure Survey for both 2009 and 2010 are in the table above. The base year is 2009.
    1. What and how much is in the CPI market basket?
  1. What did the CPI market basket cost in 2009? What was the CPI in 2009?

  1. What did the CPI market basket cost in 2010? What was the CPI in 2010?
  1. What was the inflation rate between 2009 and 2010?
Answer #1

CPI market basket cost in 2009 = 20*1+30*1+10*10 = 150

CPI in 2009 = 150/150*100 = 100

CPI market basket cost in 2010 = 15*1+45*0.75+8*11=136.75

CPI in 2010 = 136.75/150*100 = 91.16

Inflation rate = 91.16-100/100*100 = -8.84%